# SPARKE: Scalable Prompt-Aware Diversity Guidance in Diffusion Models via RKE Score

### v1

@@ A.1 Denoising Diffusion Probabilistic Models (DDPMs)
Denoising diffusion probabilistic models (DDPMs) [1, 2] define a generative process by reversing a fixed Markovian forward diffusion that progressively adds Gaussian noise to data. Given a data sample $\bm{x}_{0}\sim p_{0}(\bm{x}_{0})$, a time step $t\in[T]\triangleq\{1,\dots,T\}$, the forward process gradually adds noise to construct a noisy data point $\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{0}+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathit{\bm{I}})$ is standard Gaussian noise and $\alpha_{t}\in[0,1]$ monotonically decreases with time step $t$ to control the noise level. The diffusion model $\bm{\epsilon}_{\theta}:\mathcal{X}\times[T]\mapsto\mathcal{X}$ is trained to predict the noise $\bm{\epsilon}_{t}$ at each time step $t$, also learn the score of $p_{t}(\bm{x_{t}})$ [17, 3]:
$\min_{\theta}\mathbb{E}_{\bm{x}_{t},\bm{\epsilon}_{t},t}\left[\left\|\bm{\epsilon}_{\theta}(\bm{x}_{t},t)-\bm{\epsilon}_{t}\right\|_{2}^{2}\right]=\min_{\theta}\mathbb{E}_{\bm{x}_{t},\bm{\epsilon}_{t},t}\left[\left\|\bm{\epsilon}_{\theta}(\bm{x}_{t},t)+\sqrt{1-\alpha_{t}}\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})\right\|_{2}^{2}\right],$ (13)
The reverse process obtained by $\bm{x}_{t-1}\sim p_{t-1|t}(\bm{x}_{t-1}|\bm{x}_{t})$ is not directly computable in practice, multiple efficient reverse samplers are proposed [60, 65, 66]. In the commonly-used DDIM [60] sampler, we sample $\bm{x}_{t-1}$ by:
$\bm{x}_{t-1}=\sqrt{\alpha_{t-1}}\bm{\tilde{x}}_{0|t}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\frac{\bm{x}_{t}-\sqrt{\alpha_{t}}\bm{\tilde{x}}_{0|t}}{\sqrt{1-\alpha_{t}}}+\sigma_{t}\bm{\epsilon}_{t},$ (14)
where $\sigma_{t}$ is the DDIM parameter, and the clean sample $\bm{\tilde{x}}_{0|t}$ given $\bm{x}_{t}$ is estimated according to Tweedie’s Formula [67, 68]:
$\bm{\tilde{x}}_{0|t}=\frac{\bm{x}_{t}-\sqrt{1-\alpha_{t}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}.$ (15)
@@
In conditional generation tasks like text-to-image synthesis, diffusion models learn to approximate the conditional distribution $p(\bm{x}_{0}|\bm{y})$ given conditions $\bm{y}$ (e.g., text prompts). From the view of score functions [17, 3], we denote the conditional score as:
$\underbrace{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t}|\bm{y})}_{\text{Conditional Score}}=\underbrace{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{x}_{t})}_{\text{Unconditional Score}}+\underbrace{\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t})}_{\text{Guidance Score}},$ (16)
@@
Training-based Guidance. Training-based guidance methods include several strategies. One approach, Classifier-Guidance, initially proposed in [3, 6], requires training an additional time-dependent classifier to estimate the guidance score $f(\bm{x}_{t},t)\triangleq\mathbb{E}_{\bm{x}_{0}\sim p_{0|t}(\cdot|\bm{x}_{t})}f(\bm{x}_{0})\approx f(\bm{x}_{0})$. Alternatively, other training-based techniques involve few-shot fine-tuning of base models or the use of adapters [4, 30] to achieve conditional control.
A distinct yet training-based approach is Classifier-Free Guidance (CFG), introduced by [46]. Unlike methods requiring a separate classifier, CFG integrates the condition $\bm{y}$ as a direct input to the conditional denoising network $\bm{\epsilon}_{\theta}(\bm{x}_{t},t,\bm{y})$. It is enabled by a joint training procedure where the model also learns to make unconditional predictions by randomly dropping the condition $\bm{y}$ with a specific probability during each training iteration. During inference, CFG estimates the conditional noise prediction as follows:
$\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,\bm{y})=(1+w)\cdot\bm{\epsilon}_{\theta}(\bm{x}_{t},t,\bm{y})-w\cdot\bm{\epsilon}_{\theta}(\bm{x}_{t},t),$ (17)
@@
Training-free Guidance. An alternative category for conditional guidance is training-free guidance. Instead of requiring additional training, these methods directly introduce a time-independent conditional predictor $f$ on the estimated clean data $f(\bm{\tilde{x}}_{0|t},\bm{y})$, which can be a classifier, loss function, or energy function quantifying the alignment of a generated sample with the target condition [32, 33, 34, 69]. Then the estimation of the guidance score via training-free approaches is:
$\nabla_{\bm{x}_{t}}\log p_{t}(\bm{y}|\bm{x}_{t}):=-\nabla_{\bm{x}_{t}}f\left(\bm{\tilde{x}}_{0|t},\mathbf{y}\right),$ (18)
where the clean sample for the predictor is estimated by Tweedie’s Formula [67, 68]:
$\bm{\tilde{x}}_{0|t}=\frac{\bm{x}_{t}-\sqrt{1-\alpha_{t}}\bm{\epsilon}_{\theta}(\bm{x}_{t},t)}{\sqrt{\alpha_{t}}}.$ (19)
@@ B.1 Proof of Proposition 1
Let $K_{\mathcal{Z}}$ be the kernel matrix with entries $(K_{Z})_{ij}=k(\bm{z}^{(i)},\bm{z}^{(j)})$. The RKE loss is given by:
$\mathcal{L}_{\text{RKE}}(\bm{z}^{(1)},\ldots,\bm{z}^{(n)})=\frac{1}{\|\widetilde{K}_{Z}\|_{F}^{2}}=\frac{\mathrm{Tr}(K_{Z})^{2}}{\sum_{i=1}^{n}\sum_{j=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}},$ (20)
@@
$\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})=\frac{1}{\mathcal{L}_{\text{RKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}}{\mathrm{Tr}(K_{Z})^{2}},$ (21)
@@
Given that the kernel function $k$ is normalized, i.e., $k(\bm{z},\bm{z})=1$ for all $\bm{z}\in Z$, the trace of the kernel matrix $K_{Z}$ is:
$\mathrm{Tr}(K_{Z})=\sum_{i=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(i)})=\sum_{i=1}^{n}1=n,$
@@
Substituting this into the expression for $\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})$ in Eq. equation 21:
$\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2},$ (22)
@@
Now, we want to compute the gradient of $\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})$ with respect to $\bm{z}^{(n)}$:
@@
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})=\nabla_{\bm{z}^{(n)}}\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}\right),$ (23)
Since $1/n^{2}$ is a constant with respect to $\bm{z}^{(n)}$:
@@
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})=\frac{1}{n^{2}}\nabla_{\bm{z}^{(n)}}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}\right),$ (24)
@@
Let $S=\sum_{i=1}^{n}\sum_{j=1}^{n}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}$. We need to find $\nabla_{\bm{z}^{(n)}}S$. The terms in the sum $S$ that involve $\bm{z}^{(n)}$ are those where $i=n$ or $j=n$ (or both). We can split the sum into:
$S$ $=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}\quad(\text{Terms not involving }\bm{z}^{(n)})$
$+\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})^{2}\quad(j=n,i\neq n)$
$+\sum_{j=1}^{n-1}k(\bm{z}^{(n)},\bm{z}^{(j)})^{2}\quad(i=n,j\neq n)$
$+k(\bm{z}^{(n)},\bm{z}^{(n)})^{2},\quad(i=n,j=n)$
@@
Using the symmetry of the kernel, $k(\bm{z}^{(n)},\bm{z}^{(j)})=k(\bm{z}^{(j)},\bm{z}^{(n)})$, the second and third sums are identical:
$S=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}+2\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})^{2}+k(\bm{z}^{(n)},\bm{z}^{(n)})^{2},$ (25)
Now, we compute the gradient of $S$ with respect to $\bm{z}^{(n)}$:
$\nabla_{\bm{z}^{(n)}}S=\nabla_{\bm{z}^{(n)}}\left(\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(j)})^{2}\right)+\nabla_{\bm{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})^{2}\right)+\nabla_{\bm{z}^{(n)}}\left(k(\bm{z}^{(n)},\bm{z}^{(n)})^{2}\right),$ (26)
@@
The first term is zero because it does not depend on $\bm{z}^{(n)}$. For the second term:
$\nabla_{\bm{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})^{2}\right)$ $=2\sum_{i=1}^{n-1}\nabla_{\bm{z}^{(n)}}\left(k(\bm{z}^{(i)},\bm{z}^{(n)})^{2}\right)$ (27)
$=2\sum_{i=1}^{n-1}2k(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k(\bm{z}^{(i)},\bm{z}^{(n)})$
$=4\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k(\bm{z}^{(i)},\bm{z}^{(n)}),$
@@
For the third term, since $k(\bm{z},\bm{z})=1$ (normalized kernel), this term is constant:
$\nabla_{\bm{z}^{(n)}}\left(k(\bm{z}^{(n)},\bm{z}^{(n)})^{2}\right)=\nabla_{\bm{z}^{(n)}}(1^{2})=\nabla_{\bm{z}^{(n)}}(1)=\bm{0},$ (28)
According to Eq. equation 27 and Eq. equation 28, we have the gradeitn for $S$:
$\nabla_{\bm{z}^{(n)}}S=4\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k(\bm{z}^{(i)},\bm{z}^{(n)}),$ (29)
@@
Substituting this back into the gradient of $\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})$ in Eq. equation 24:
@@
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{IRKE}}(\bm{z}^{(1)},\dots,\bm{z}^{(n)})$ $=\frac{1}{n^{2}}\left(4\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k(\bm{z}^{(i)},\bm{z}^{(n)})\right)$ (30)
$=\frac{4}{n^{2}}\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k(\bm{z}^{(i)},\bm{z}^{(n)}),$
@@
$\nabla_{\bm{z}^{(n)}}\mathcal{L}_{\text{IRKE}}\propto\sum_{i=1}^{n-1}k(\bm{z}^{(i)},\bm{z}^{(n)})\nabla_{\bm{z}^{(n)}}k(\bm{z}^{(i)},\bm{z}^{(n)}).$ (31)

### v2

@@ 6 Conclusion and Limitations
Acknowledgments
The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. The work of Amin Gohari is supported by CUHK Direct Research Grants with CUHK Project No. 4055270. Also, the authors would like to sincerely thank the anonymous reviewers for their insightful feedback and constructive suggestions.
@@ A.1 Denoising Diffusion Probabilistic Models (DDPMs)
Denoising diffusion probabilistic models (DDPMs) sohl2015deep ; ho2020denoising define a generative process by reversing a fixed Markovian forward diffusion that progressively adds Gaussian noise to data. Given a data sample $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x}_{0})$, a time step $t\in[T]\triangleq\{1,\dots,T\}$, the forward process gradually adds noise to construct a noisy data point $\boldsymbol{x}_{t}=\sqrt{\alpha_{t}}\boldsymbol{x}_{0}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{t}$, where $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathit{\boldsymbol{I}})$ is standard Gaussian noise and $\alpha_{t}\in[0,1]$ monotonically decreases with time step $t$ to control the noise level. The diffusion model $\boldsymbol{\epsilon}_{\theta}:\mathcal{X}\times[T]\mapsto\mathcal{X}$ is trained to predict the noise $\boldsymbol{\epsilon}_{t}$ at each time step $t$, also learn the score of $p_{t}(\boldsymbol{x_{t}})$ song2019generative ; song2020score :
$\min_{\theta}\mathbb{E}_{\boldsymbol{x}_{t},\boldsymbol{\epsilon}_{t},t}\left[\left\|\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)-\boldsymbol{\epsilon}_{t}\right\|_{2}^{2}\right]=\min_{\theta}\mathbb{E}_{\boldsymbol{x}_{t},\boldsymbol{\epsilon}_{t},t}\left[\left\|\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)+\sqrt{1-\alpha_{t}}\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})\right\|_{2}^{2}\right],$ (12)
The reverse process obtained by $\boldsymbol{x}_{t-1}\sim p_{t-1|t}(\boldsymbol{x}_{t-1}|\boldsymbol{x}_{t})$ is not directly computable in practice, multiple efficient reverse samplers are proposed song2020denoising ; lu2022dpm ; karras2022elucidating . In the commonly-used DDIM song2020denoising sampler, we sample $\boldsymbol{x}_{t-1}$ by:
$\boldsymbol{x}_{t-1}=\sqrt{\alpha_{t-1}}\boldsymbol{\tilde{x}}_{0|t}+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\frac{\boldsymbol{x}_{t}-\sqrt{\alpha_{t}}\boldsymbol{\tilde{x}}_{0|t}}{\sqrt{1-\alpha_{t}}}+\sigma_{t}\boldsymbol{\epsilon}_{t},$ (13)
where $\sigma_{t}$ is the DDIM parameter, and the clean sample $\boldsymbol{\tilde{x}}_{0|t}$ given $\boldsymbol{x}_{t}$ is estimated according to Tweedie’s Formula efron2011tweedie ; robbins1992empirical :
$\boldsymbol{\tilde{x}}_{0|t}=\frac{\boldsymbol{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)}{\sqrt{\alpha_{t}}}.$ (14)
@@
In conditional generation tasks like text-to-image synthesis, diffusion models learn to approximate the conditional distribution $p(\boldsymbol{x}_{0}|\boldsymbol{y})$ given conditions $\boldsymbol{y}$ (e.g., text prompts). From the view of score functions song2019generative ; song2020score , we denote the conditional score as:
$\underbrace{\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t}|\boldsymbol{y})}_{\text{Conditional Score}}=\underbrace{\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{x}_{t})}_{\text{Unconditional Score}}+\underbrace{\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{y}|\boldsymbol{x}_{t})}_{\text{Guidance Score}},$ (15)
@@
Training-based Guidance. Training-based guidance methods include several strategies. One approach, Classifier-Guidance, initially proposed in song2020score ; dhariwal2021diffusion , requires training an additional time-dependent classifier to estimate the guidance score $f(\boldsymbol{x}_{t},t)\triangleq\mathbb{E}_{\boldsymbol{x}_{0}\sim p_{0|t}(\cdot|\boldsymbol{x}_{t})}f(\boldsymbol{x}_{0})\approx f(\boldsymbol{x}_{0})$. Alternatively, other training-based techniques involve few-shot fine-tuning of base models or the use of adapters rombach2022high ; ruiz2023dreambooth to achieve conditional control.
A distinct yet training-based approach is Classifier-Free Guidance (CFG), introduced by ho2022classifier . Unlike methods requiring a separate classifier, CFG integrates the condition $\boldsymbol{y}$ as a direct input to the conditional denoising network $\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{y})$. It is enabled by a joint training procedure where the model also learns to make unconditional predictions by randomly dropping the condition $\boldsymbol{y}$ with a specific probability during each training iteration. During inference, CFG estimates the conditional noise prediction as follows:
$\hat{\boldsymbol{\epsilon}}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{y})=(1+w)\cdot\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t,\boldsymbol{y})-w\cdot\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t),$ (16)
@@
Training-free Guidance. An alternative category for conditional guidance is training-free guidance. Instead of requiring additional training, these methods directly introduce a time-independent conditional predictor $f$ on the estimated clean data $f(\boldsymbol{\tilde{x}}_{0|t},\boldsymbol{y})$, which can be a classifier, loss function, or energy function quantifying the alignment of a generated sample with the target condition bansal2023universal ; yu2023freedom ; ye2024tfg ; shen2024understanding . Then the estimation of the guidance score via training-free approaches is:
$\nabla_{\boldsymbol{x}_{t}}\log p_{t}(\boldsymbol{y}|\boldsymbol{x}_{t}):=-\nabla_{\boldsymbol{x}_{t}}f\left(\boldsymbol{\tilde{x}}_{0|t},\mathbf{y}\right),$ (17)
where the clean sample for the predictor is estimated by Tweedie’s Formula efron2011tweedie ; robbins1992empirical :
$\boldsymbol{\tilde{x}}_{0|t}=\frac{\boldsymbol{x}_{t}-\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{x}_{t},t)}{\sqrt{\alpha_{t}}}.$ (18)
@@ B.1 Proof of Proposition 1
Let $K_{\mathcal{Z}}$ be the kernel matrix with entries $(K_{Z})_{ij}=k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})$. The RKE loss is given by:
$\mathcal{L}_{\text{RKE}}(\boldsymbol{z}^{(1)},\ldots,\boldsymbol{z}^{(n)})=\frac{1}{\|\widetilde{K}_{Z}\|_{F}^{2}}=\frac{\mathrm{Tr}(K_{Z})^{2}}{\sum_{i=1}^{n}\sum_{j=1}^{n}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}},$ (19)
@@
$\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})=\frac{1}{\mathcal{L}_{\text{RKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})}=\frac{\sum_{i=1}^{n}\sum_{j=1}^{n}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}}{\mathrm{Tr}(K_{Z})^{2}},$ (20)
@@
Given that the kernel function $k$ is normalized, i.e., $k(\boldsymbol{z},\boldsymbol{z})=1$ for all $\boldsymbol{z}\in Z$, the trace of the kernel matrix $K_{Z}$ is:
$\mathrm{Tr}(K_{Z})=\sum_{i=1}^{n}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(i)})=\sum_{i=1}^{n}1=n,$
@@
Substituting this into the expression for $\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})$ in Eq. (20):
$\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2},$ (21)
@@
Now, we want to compute the gradient of $\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})$ with respect to $\boldsymbol{z}^{(n)}$:
@@
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})=\nabla_{\boldsymbol{z}^{(n)}}\left(\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}\right),$ (22)
Since $1/n^{2}$ is a constant with respect to $\boldsymbol{z}^{(n)}$:
@@
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})=\frac{1}{n^{2}}\nabla_{\boldsymbol{z}^{(n)}}\left(\sum_{i=1}^{n}\sum_{j=1}^{n}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}\right),$ (23)
@@
Let $S=\sum_{i=1}^{n}\sum_{j=1}^{n}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}$. We need to find $\nabla_{\boldsymbol{z}^{(n)}}S$. The terms in the sum $S$ that involve $\boldsymbol{z}^{(n)}$ are those where $i=n$ or $j=n$ (or both). We can split the sum into:
$S$ $=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}\quad(\text{Terms not involving }\boldsymbol{z}^{(n)})$
$+\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}\quad(j=n,i\neq n)$
$+\sum_{j=1}^{n-1}k(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(j)})^{2}\quad(i=n,j\neq n)$
$+k(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2},\quad(i=n,j=n)$
@@
Using the symmetry of the kernel, $k(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(j)})=k(\boldsymbol{z}^{(j)},\boldsymbol{z}^{(n)})$, the second and third sums are identical:
$S=\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}+2\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}+k(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2},$ (24)
Now, we compute the gradient of $S$ with respect to $\boldsymbol{z}^{(n)}$:
$\nabla_{\boldsymbol{z}^{(n)}}S=\nabla_{\boldsymbol{z}^{(n)}}\left(\sum_{i=1}^{n-1}\sum_{j=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(j)})^{2}\right)+\nabla_{\boldsymbol{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}\right)+\nabla_{\boldsymbol{z}^{(n)}}\left(k(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2}\right),$ (25)
@@
The first term is zero because it does not depend on $\boldsymbol{z}^{(n)}$. For the second term:
$\nabla_{\boldsymbol{z}^{(n)}}\left(2\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}\right)$ $=2\sum_{i=1}^{n-1}\nabla_{\boldsymbol{z}^{(n)}}\left(k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})^{2}\right)$ (26)
$=2\sum_{i=1}^{n-1}2k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})$
$=4\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}),$
@@
For the third term, since $k(\boldsymbol{z},\boldsymbol{z})=1$ (normalized kernel), this term is constant:
$\nabla_{\boldsymbol{z}^{(n)}}\left(k(\boldsymbol{z}^{(n)},\boldsymbol{z}^{(n)})^{2}\right)=\nabla_{\boldsymbol{z}^{(n)}}(1^{2})=\nabla_{\boldsymbol{z}^{(n)}}(1)=\boldsymbol{0},$ (27)
According to Eq. (26) and Eq. (27), we have the gradeitn for $S$:
$\nabla_{\boldsymbol{z}^{(n)}}S=4\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}),$ (28)
@@
Substituting this back into the gradient of $\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})$ in Eq. (23):
@@
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{IRKE}}(\boldsymbol{z}^{(1)},\dots,\boldsymbol{z}^{(n)})$ $=\frac{1}{n^{2}}\left(4\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\right)$ (29)
$=\frac{4}{n^{2}}\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}),$
@@
$\nabla_{\boldsymbol{z}^{(n)}}\mathcal{L}_{\text{IRKE}}\propto\sum_{i=1}^{n-1}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)})\nabla_{\boldsymbol{z}^{(n)}}k(\boldsymbol{z}^{(i)},\boldsymbol{z}^{(n)}).$ (30)